Reports 1-1 of 1 Clear search Modify search
AdV-SIN (Pre-commissioning)
polini - 22:21 Friday 10 September 2021 (53126) Print this report
Characterization of new PSD responses on SQB1

Today, I characterized the responses of the two PSDs (logbook entry #52701) we replaced on SQB1.

These PSDs are the PSD1 and the PSD3: PSD1 is the one reading the beam arriving from EQB1, in  reflection from the steering mirror with picomotors SQB1_GM12 and PSD3 is the one in reflection from the filter cavity.

First of all, I aligned with the picomotors the green beam on PSD1. The starting values on the PSDs is reported in the attached plot. PSD2 is not well aligned but since we didn't replace it, we already have the sensing coefficients for this one. We are planning to put picomotors in front of this PSD in order to be able to align it remotely.

I turned off the driflt control of the cavity otherwise the incoming green beam was moving during the scans of SQB1.I re-aligned the cavity in the usual way: moving the IM and EM maximizing the flashes. After turning off the drift control, the peaks were or ~1V instead of 7V. To re-align the cavity wasn't trivial. After several tries I got peaks of 6V (attached plot).

Once in this situation, I was able to perform the scans of the angles and positions of SQB1.

The following scans have been performed:

  • TX: 2021-09-10-18h41m12-UTC    info polini       'Bench Set points:LC_TX_set [LC_TX=300,ramp time=100]' sent to SQB1_LC
  • TY: 2021-09-10-17h59m20-UTC    info polini       'Bench Set points:LC_TY_set [LC_TY=-250,ramp time=100]' sent to SQB1_LC
  • TZ: 2021-09-10-18h04m37-UTC    info polini       'Bench Set points:LC_TZ_set [LC_TZ=300,ramp time=100]' sent to SQB1_LC
  • X: 2021-09-10-18h13m27-UTC    info polini       'Bench Set points:LC_X_set [LC_X=-800,ramp time=130]' sent to SQB1_LC
  • Y: 2021-09-10-18h24m22-UTC    info polini       'Bench Set points:LC_Y_set [LC_Y=100,ramp time=100]' sent to SQB1_LC
  • Z: 2021-09-10-18h35m15-UTC    info polini       'Bench Set points:LC_TZ_set [LC_Z=510,ramp time=100]' sent to SQB1_LC
        2021-09-10-18h37m20-UTC    info polini       'Bench Set points:LC_TZ_set [LC_Z=510,ramp time=100]' sent to SQB1_LC

At the end of the scans, I engaged the drift control of the filter cavity back and the cavity is back locked at 7.1V (see last picture attached).

Images attached to this report
Search Help
×

Warning

×